# A bound on the quantum value of all compiled nonlocal games

### v1

@@ 4.3. Asymptotic security for any noncommutative polynomial
(4.5) $\sigma^{\lambda}_{x}=\sum_{a\in\mathcal{O}_{A}}\sigma^{\lambda}_{xa}=\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\xi\leftarrow\mathsf{Enc}(\mathsf{sk},x)}\sum_{\alpha\in\mathsf{CT}}A_{\xi\alpha}^{\lambda}\lvert\psi^{\lambda}\rangle\langle\mathopen{}\psi^{\lambda}\rvert(A_{\xi\alpha}^{\lambda})^{*}.$
@@
Proposition 4.6 is a generalization of [CMM+24, Lem. 21] and is proved in a similar fashion, by using block encodings.
@@
$tH=\left(\langle 0\rvert^{\otimes m}\otimes\mathbbm{1}\right)U\left(\lvert 0\rangle^{\otimes m}\otimes\mathbbm{1}\right),\quad\text{i.e.}\quad U=\begin{pmatrix}tH&*\\ *&*\end{pmatrix}$
@@
We first note that for any fixed $y\in\mathcal{I}_{B}$ and $b\in\mathcal{O}_{B}$, the POVM elements $\{B^{\lambda}_{yb}\}$ have a natural QPT block encoding. This follows from equation (4.2), by the same reasoning as in [GSLW19, Lem. 26], which also shows that the resulting block encodings have parameter $t=1$ and $m=N$, where $N$ denotes the total number of bits in the binary representation of $y$ and $b$.
Now suppose that $P$ is a monomial of degree $D$ and let $M_{\lambda}=P(\{B^{\lambda}_{yb}\})$. It follows from [GSLW19, Lem. 30] that $\{M_{\lambda}\}$ admits a QPT block encoding with $t=1$ and $m=DN$, simply by concatenating the individual block encodings in a suitable way (cf. [CMM+24, Lem 2.18]). Even though each POVM element is an observable, the operators $M_{\lambda}$ need not be Hermitian, so we cannot apply Lemma 4.8 directly. Instead, we observe that it suffices to prove the claim for the observables $\operatorname{Re}(M_{\lambda})=(M_{\lambda}+M_{\lambda}^{*})/2$ and $\operatorname{Im}(M_{\lambda})=(M_{\lambda}-M_{\lambda}^{*})/(2\imath)$. To this end, we first note that $\{M_{\lambda}^{*}\}$ also admits a QPT block encoding with the same parameters as $\{M_{\lambda}\}$, as the former family corresponds to the monomial obtained by reversing $P$. Then it follows from [GSLW19, Lem. 29], by taking the controlled unitaries corresponding to these QPT block encodings, along with fixed state-preparation pairs for the two desired linear combinations, that the two families $\{\operatorname{Re}(M_{\lambda})\}$ and $\{\operatorname{Im}(M_{\lambda})\}$ admit QPT block encodings (cf. [CMM+24, Lem 2.17]). Now the claim follows from Lemma 4.8, the triangle inequality, and the fact that nonnegative linear combinations of negligible functions are negligible. ∎
5. Sequential characterizations of nonlocal correlations
Motivated by the two-round structure of a compiled game, we consider sequential games and strategies. Without further constraints, the resulting correlations can even be signaling, but we identify an information-theoretic property motivated by Proposition 4.6 that ensures that the resulting correlations are exactly quantum (in the finite-dimensional case) respectively commuting operator correlations (in the infinite-dimensional case). We also give a classical version of this result.
5.1. Sequential games
@@
A sequential game can also be interpreted as a two-player game where the first player can pass some information (depending on their question) to the second player before the latter has to respond with their answer.
@@
5.2. Classical strategies and correlations
While our main interest are quantum strategies, we first discuss the classical case to build some intuition.
@@ 5.2. Classical strategies and correlations
    probability distributions $\{p(a,\omega|x):x\in\mathcal{I}_{A}\}$ with outcomes in $\mathcal{I}_{A}\times\Omega$, where $\Omega$ is a (without loss of generality) finite set,
@@
Consider a classical strategy for $\mathcal{G}_{\textnormal{seq}}$ and suppose that the distributions $p(\omega|x)=\sum_{a\in\mathcal{O}_{A}}p(a,\omega|x)$ are the same for all $x\in\mathcal{I}_{A}$. Then the resulting correlation is a nonlocal classical correlation, that is, in $C_{c}$.
@@
Conversely, any classical strategy for the nonlocal game (Definition 3.3) gives rise to one for the sequential game that satisfies the hypotheses of Proposition 5.4. Simply set $p(a,\omega|x):=\gamma(\omega)p_{\omega}(a|x)$ and use the same $q_{\omega}(b|y)$ as in the nonlocal strategy.
@@
5.3. Quantum strategies and correlations
@@
Such a quantum strategy gives rise to a correlation
@@
Let $\mathcal{H}$ be a Hilbert space, $\rho$ be a state on $\mathcal{H}$, and $\{\Phi_{xa}\}_{x\in\mathcal{I}_{A},a\in\mathcal{O}_{A}}$ be a collection of completely positive maps such that $\sum_{a\in\mathcal{O}_{A}}\Phi_{xa}$ is trace-preserving for every $x\in\mathcal{I}_{A}$. [^6] A collection of completely positive maps $\{\Psi_{a}\}_{a\in\mathcal{O}}$ such that $\sum_{a\in\mathcal{O}_{A}}\Psi_{xa}$ is trace-preserving is called a quantum instrument. It describes the most general quantum evolution that has a classical outcome (measurement result) as well as a quantum one (post-measurement state) [Wil17, §4.6.8]. Then the operators $\sigma_{xa}=\Phi_{xa}(\rho)$ satisfy the assumptions in ii of Definition 5.7. Conversely, any collection of operators as in ii arises in this way.
@@
A quantum strategy for $\mathcal{G}_{\textnormal{seq}}$ is strongly non-signaling if there exists a $C^{*}$-algebra $\mathscr{B}\subseteq\mathcal{B}(\mathcal{H})$ containing the operators $\{B_{yb}\}_{y\in\mathcal{I}_{B},b\in\mathcal{O}_{B}}$ such that the following condition holds: for all $x,x^{\prime}\in\mathcal{I}_{A}$ and for all $B\in\mathscr{B}$, we have
(5.1) $\operatorname{tr}(\sigma_{x}B)=\operatorname{tr}(\sigma_{x^{\prime}}B).$
We can always take $\mathscr{B}$ to be the $C^{*}$-algebra generated by the operators $\{B_{yb}\}$, because it is contained in any other $C^{*}$-algebra that contains these elements. Moreover, by continuity it suffices to verify (5.1) on the dense set of noncommutative polynomials in these operators. [^7] However, it does not suffice to only require that (5.1) holds for the generators $B\in\{B_{yb}\}_{y\in\mathcal{I}_{B},b\in\mathcal{O}_{B}}$. We record this useful observation.
@@
(5.2) $\operatorname{tr}\bigl{(}\sigma_{x}\,P(\{B_{yb}\})\bigr{)}=\operatorname{tr}\bigl{(}\sigma_{x^{\prime}}\,P(\{B_{yb}\})\bigr{)}.$
@@
However, to build intuition we first consider finite-dimensional quantum sequential strategies. In this case we can prove that the strong non-signaling property ensures that these give rise to nonlocal quantum correlations in the sense of Definition 3.4.
Proposition 5.11.
Consider a quantum strategy for $\mathcal{G}_{\textnormal{seq}}$ that satisfies the strong non-signaling condition. If $\dim(\mathcal{H})<\infty$, then the resulting correlation is a nonlocal quantum correlation, that is, in $C_{q}$.
@@
As explained in Remark 5.17 we may assume that $\mathscr{B}=\mathcal{B}(\mathcal{H})$ and hence $\sigma_{x}=\sigma_{x^{\prime}}$ for all $x,x^{\prime}\in\mathcal{I}_{A}$, and therefore we let $\sigma=\sigma_{x}$. Let $\mathcal{H}_{A}:=\mathbb{C}^{\mathcal{O}_{A}}\otimes\mathcal{H}_{A^{\prime}}$ and $\mathcal{H}_{A^{\prime}}:=\mathcal{H}_{B}:=\mathcal{H}$. We think of the operators $\sigma_{xa}$ as acting on $\mathcal{H}_{B}$ and choose purifications $\lvert\psi_{xa}\rangle\in\mathcal{H}_{A^{\prime}}\otimes\mathcal{H}_{B}$ for each $x\in\mathcal{I}_{A}$ and $a\in\mathcal{O}_{A}$. Then the states
@@
which shows that $p(a,b|x,y)$ is a quantum correlation, with Hilbert spaces $\mathcal{H}_{A},\mathcal{H}_{B}$, initial state $\lvert\psi_{x_{0}}\rangle$, and POVM elements $A_{xa}:=U_{xx_{0}}^{*}P_{a}U_{xx_{0}}$ and $B_{yb}$ for all $a\in\mathcal{O}_{A},b\in\mathcal{O}_{B},x\in\mathcal{I}_{A}$, and $y\in\mathcal{I}_{B}$. ∎
Conversely, any quantum strategy for the nonlocal game (Definition 3.4) gives rise to one for the sequential game that satisfies the hypotheses of Proposition 5.11. Simply let $\mathcal{H}=\mathcal{H}_{B}$, $\sigma_{xa}=\operatorname{tr}_{A}((M_{xa}\otimes\mathbbm{1}_{B})\lvert\psi\rangle\langle\mathopen{}\psi\rvert)$, and use the operators $B_{yb}=N_{yb}$. Then strong non-signaling is satisfied, e.g., for $\mathscr{B}=\mathcal{B}(\mathcal{H})$.
@@
The quantum correlation set $C_{q}$ consists precisely of the correlations produced by finite-dimensional quantum sequential strategies satisfying the strong non-signaling condition of Definition 5.9.
If we do not require the Hilbert space to be finite dimensional, we find that the strong non-signaling condition precisely characterizes the commuting operator correlations in the sense of Definition 3.5. To establish this result, it is useful to define the following equivalent $C^{*}$-algebraic model.

### v3

@@ 4.3 Asymptotic security for any noncommutative polynomial
$\sigma^{\lambda}_{x}=\sum_{a\in\mathcal{O}_{A}}\sigma^{\lambda}_{xa}=\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\xi\leftarrow\mathsf{Enc}(\mathsf{sk},x)}\sum_{\alpha\in\mathsf{CT}}A_{\xi\alpha}^{\lambda}\lvert\psi^{\lambda}\rangle\langle\mathopen{}\psi^{\lambda}\rvert(A_{\xi\alpha}^{\lambda})^{*}.$ (4.5)
@@
Proposition 4.6 is a generalization of [15, Lem. 21] and is proved in a similar fashion, by using block encodings.
@@
$tM=\left(\langle 0\rvert^{\otimes m}\otimes\mathbbm{1}\right)U\left(\lvert 0\rangle^{\otimes m}\otimes\mathbbm{1}\right),\quad\text{i.e.}\quad U=\begin{pmatrix}tM&*\\ *&*\end{pmatrix}$
@@
We first note that for any fixed $y\in\mathcal{I}_{B}$ and $b\in\mathcal{O}_{B}$, the POVM elements $\{B^{\lambda}_{yb}\}$ have a natural QPT block encoding. This follows from equation (4.2), by the same reasoning as in [19, Lem. 26], which also shows that the resulting block encodings have parameter $t=1$ and $m=N$, where $N$ denotes the total number of bits in the binary representation of $y$ and $b$.
Now suppose that $P$ is a monomial of degree $D$ and let $M_{\lambda}=P(\{B^{\lambda}_{yb}\})$. It follows from [19, Lem. 30] that $\{M_{\lambda}\}$ admits a QPT block encoding with $t=1$ and $m=DN$, simply by concatenating the individual block encodings in a suitable way (cf. [15, Lem 2.18]). Even though each POVM element is an observable, the operators $M_{\lambda}$ need not be Hermitian, so we cannot apply Lemma 4.8 directly. Instead, we observe that it suffices to prove the claim for the observables $\operatorname{Re}(M_{\lambda})=(M_{\lambda}+M_{\lambda}^{*})/2$ and $\operatorname{Im}(M_{\lambda})=(M_{\lambda}-M_{\lambda}^{*})/(2\imath)$. To this end, we first note that $\{M_{\lambda}^{*}\}$ also admits a QPT block encoding with the same parameters as $\{M_{\lambda}\}$, as the former family corresponds to the monomial obtained by reversing $P$. Then it follows from [19, Lem. 29], by taking the controlled unitaries corresponding to these QPT block encodings, along with fixed state-preparation pairs for the two desired linear combinations, that the two families $\{\operatorname{Re}(M_{\lambda})\}$ and $\{\operatorname{Im}(M_{\lambda})\}$ admit QPT block encodings (cf. [15, Lem 2.17]). Now the claim follows from Lemma 4.8, the triangle inequality, and the fact that nonnegative linear combinations of negligible functions are negligible. ∎
5 Sequential characterizations of nonlocal correlations
Motivated by the two-round structure of a compiled game, we consider sequential games and strategies. Without further constraints, the resulting correlations can even be signaling, but we find that a natural information-theoretic property motivated by Proposition 4.6 ensures that the resulting correlations are nonlocal ones. We describe the sequential setting in Section 5.1. In Sections 5.2 and 5.3 we discuss how to characterize classical as well as (finite-dimensional and spatial) quantum correlations in this setting. In Section 5.4 we prove the main result of this section: a “sequential” characterization of commuting operator correlations (Theorem 5.15), which to the best of our knowledge has not appeared in the literature before.
Only the latter is required for the main results of this article, and we invite readers interested only in those to proceed directly to the self-contained Section 5.4. The pedagogical Sections 5.2 and 5.3 discuss more concrete settings that allow to gain intuition for the more abstract algebraic results of Section 5.4, and the results therein can be obtained from prior work, as we explain below.
5.1 Sequential games
@@
A sequential game can also be interpreted as a two-player game where the first player can pass some information (depending on their question) to the second player before the latter has to respond with their answer. Sequential games have been investigated from different perspectives; see, e.g., [7] and the references therein.
We note that this setting can also be translated into the language of prepare-and-measure scenarios studied in the contextuality literature. In particular, [46] proposes a general map between nonlocal scenarios and strategies on the one hand and certain prepare-and-measure contextuality scenarios and preparations on the other hand (their introduction also gives an account of the motivations of this line of research and prior works). Corollaries 5.5 and 5.13 can be obtained from their work (which in turn builds on [34]), while our Theorems 5.15 and 5.17 resolve an open question left in their work.
@@
5.2 Classical strategies and correlations
While our main interest is quantum strategies, we first discuss the classical case to build some intuition.
@@ 5.2 Classical strategies and correlations
    probability distributions $\{p(a,\omega|x):x\in\mathcal{I}_{A}\}$ with outcomes in $\mathcal{O}_{A}\times\Omega$, where $\Omega$ is a (without loss of generality) finite set,
@@
Consider a classical strategy for $\mathcal{G}_{\textnormal{seq}}$ and suppose that the distributions $p(\omega|x)=\sum_{a\in\mathcal{O}_{A}}p(a,\omega|x)$ are the same for all $x\in\mathcal{I}_{A}$. Then the resulting correlation is a (nonlocal) classical correlation, that is, in $C_{c}$.
@@
Conversely, any classical strategy for the nonlocal game (Definition 3.3) gives rise to one for the sequential game that satisfies the hypotheses of Proposition 5.4. Simply set $p(a,\omega|x):=\gamma(\omega)p_{\omega}(a|x)$ and use the same $q_{\omega}(b|y)$ as in the nonlocal strategy. We thus obtain the following characterization, which can also be obtained by translating [46, App. D] into the language of sequential games:
@@
5.3 Quantum strategies and correlations
@@
(We say that the strategy is finite-dimensional if the Hilbert space $\mathcal{H}$ is finite-dimensional.) Such a quantum strategy gives rise to a correlation
@@
Let $\mathcal{H}$ be a Hilbert space, $\rho$ be a state on $\mathcal{H}$, and $\{\Phi_{xa}\}_{x\in\mathcal{I}_{A},a\in\mathcal{O}_{A}}$ be a collection of completely positive maps such that $\sum_{a\in\mathcal{O}_{A}}\Phi_{xa}$ is trace-preserving for every $x\in\mathcal{I}_{A}$. [^6] A collection of completely positive maps $\{\Psi_{a}\}_{a\in\mathcal{O}}$ such that $\sum_{a\in\mathcal{O}_{A}}\Psi_{xa}$ is trace-preserving is called a quantum instrument. It describes the most general quantum evolution that has a classical outcome (measurement result) as well as a quantum one (post-measurement state) [45, §4.6.8]. Then the operators $\sigma_{xa}=\Phi_{xa}(\rho)$ satisfy the assumptions in ii of Definition 5.7. Conversely, any collection of operators as in ii arises in this way.
@@
We say that a quantum strategy is strongly non-signaling if there exists a $C^{*}$-algebra $\mathscr{B}\subseteq\mathcal{B}(\mathcal{H})$ containing the operators $\{B_{yb}\}_{y\in\mathcal{I}_{B},b\in\mathcal{O}_{B}}$ such the following condition holds: for all $x,x^{\prime}\in\mathcal{I}_{A}$ and for all $B\in\mathscr{B}$, we have
$\operatorname{tr}(\sigma_{x}B)=\operatorname{tr}(\sigma_{x^{\prime}}B).$ (5.1)
We say that $S$ is strongly non-signaling with respect to $\mathscr{B}$ to indicate $\mathscr{B}$ explicitly.
Note that a quantum strategy is strongly non-signaling if, and only if, it is so with respect to the $C^{*}$-algebra generated by the operators $\{B_{yb}\}$ because it is contained in any other $C^{*}$-algebra that contains these elements. By continuity it suffices to verify (5.1) on the dense set of noncommutative polynomials in these operators. [^7] However, it does not suffice to only require that (5.1) holds for the generators $B\in\{B_{yb}\}_{y\in\mathcal{I}_{B},b\in\mathcal{O}_{B}}$. We record this useful observation.
@@
$\operatorname{tr}\bigl{(}\sigma_{x}\,P(\{B_{yb}\})\bigr{)}=\operatorname{tr}\bigl{(}\sigma_{x^{\prime}}\,P(\{B_{yb}\})\bigr{)}.$ (5.2)
@@
To build intuition we first consider the special case of quantum sequential strategies that are strongly non-signaling with respect to $\mathcal{B}(\mathcal{H})$ for some Hilbert space $\mathcal{H}$. In other words, we consider the situation that $\sigma_{x}=\sigma_{x^{\prime}}$ for all $x,x^{\prime}\in\mathcal{I}_{A}$. This situation was also studied in [34] in a different context. There, the corresponding correlations were called “quansal” and it was proved that every quansal correlation is in $C_{qs}$. We give a self-contained proof of this result:
Proposition 5.11 ([34, Lem. 4]).
Consider a quantum strategy for $\mathcal{G}_{\textnormal{seq}}$ that is strongly non-signaling with respect to $\mathcal{B}(\mathcal{H})$. Then the resulting correlation is a (nonlocal) spatial quantum correlation, that is, in $C_{qs}$.
@@
Since the quantum strategy is strong non-signaling condition with respect to $\mathcal{B}(\mathcal{H})$, it holds $\sigma_{x}=\sigma_{x^{\prime}}$ for all $x,x^{\prime}\in\mathcal{I}_{A}$, and therefore we let $\sigma=\sigma_{x}$. Let $\mathcal{H}_{A}:=\mathbb{C}^{\mathcal{O}_{A}}\otimes\mathcal{H}_{A^{\prime}}$ and $\mathcal{H}_{A^{\prime}}:=\mathcal{H}_{B}:=\mathcal{H}$. We think of the operators $\sigma_{xa}$ as acting on $\mathcal{H}_{B}$ and choose purifications $\lvert\psi_{xa}\rangle\in\mathcal{H}_{A^{\prime}}\otimes\mathcal{H}_{B}$ for each $x\in\mathcal{I}_{A}$ and $a\in\mathcal{O}_{A}$. Then the states
@@
which shows that $p(a,b|x,y)$ is a quantum spatial correlation, with Hilbert spaces $\mathcal{H}_{A}$, $\mathcal{H}_{B}$, initial state $\lvert\psi_{x_{0}}\rangle$, and POVM elements $A_{xa}:=U_{xx_{0}}^{*}P_{a}U_{xx_{0}}$ and $B_{yb}$ for all $a\in\mathcal{O}_{A},b\in\mathcal{O}_{B},x\in\mathcal{I}_{A}$, and $y\in\mathcal{I}_{B}$. ∎
It is clear by inspection of the preceding proof that if $\mathcal{H}$ is finite-dimensional then the resulting correlation is a quantum correlation, that is, in $C_{q}$. In fact, in the finite-dimensional case, we do not have to assume that $\mathscr{B}=\mathcal{B}(\mathcal{H})$:
Proposition 5.12.
Consider a finite-dimensional quantum strategy for $\mathcal{G}_{\textnormal{seq}}$ that is strongly non-signaling. Then the resulting correlation is a (nonlocal) quantum correlation, that is, in $C_{q}$.
Proof.
By the preceding discussion, it suffices to show that for every such strategy one can find some other strategy that produces the same correlations but is strongly non-signaling for $\mathcal{B}(\mathcal{H})$. Then the result follows from Proposition 5.11. By the classification of finite-dimensional *-algebras, we may assume that
$\mathcal{H}=\bigoplus_{i}(\mathbb{C}^{n_{i}}\otimes\mathbb{C}^{m_{i}}),\;\;\mathscr{B}=\bigoplus_{i}M_{n_{i}}(\mathbb{C})\otimes\mathbbm{1}_{m_{i}},\;\;\text{and}\;\;\mathscr{B}^{\prime}=\bigoplus_{i}\mathbbm{1}_{n_{i}}\otimes M_{m_{i}}(\mathbb{C}),$
where $\mathscr{B}^{\prime}$ denotes the commutant of $\mathscr{B}$. Denote by $P_{i}$ the projection onto the $i$th direct summand and define positive operators $\tilde{\sigma}_{xa}=\bigoplus_{i}\operatorname{tr}_{M_{m_{i}}(\mathbb{C})}(P_{i}\sigma_{xa}P_{i})\otimes\frac{\mathbbm{1}_{m_{i}}}{m_{i}}$ for $x\in\mathcal{I}_{A}$ and $a\in\mathcal{O}_{A}$, where $\operatorname{tr}_{M_{m_{i}}(\mathbb{C})}$ is the partial trace over the second tensor factor of the $i$th summand. Let $B=\bigoplus_{i}B^{(i)}\otimes\mathbbm{1}_{m_{i}}$ be an arbitrary element in $\mathscr{B}$. We compute:
$\operatorname{tr}(\sigma_{xa}B)$ $=\sum_{i}\operatorname{tr}\bigl{(}\sigma_{xa}P_{i}(B^{(i)}\otimes\mathbbm{1}_{m_{i}})P_{i}\bigr{)}$
$=\sum_{i}\operatorname{tr}\bigl{(}P_{i}\sigma_{xa}P_{i}(B^{(i)}\otimes\mathbbm{1}_{m_{i}})\bigr{)}$
$=\sum_{i}\operatorname{tr}\bigl{(}\operatorname{tr}_{M_{m_{i}}(\mathbb{C})}(P_{i}\sigma_{xa}P_{i})B^{(i)}\bigr{)}$
$=\sum_{i}\operatorname{tr}\bigl{(}(\operatorname{tr}_{M_{m_{i}}(\mathbb{C})}(P_{i}\sigma_{xa}P_{i})\otimes\tfrac{\mathbbm{1}_{m_{i}}}{m_{i}})(B^{(i)}\otimes\mathbbm{1}_{m_{i}})\bigr{)}$
$=\operatorname{tr}(\tilde{\sigma}_{xa}B).$
On the one hand, this shows that the operators $\tilde{\sigma}_{xa}$ produce the same correlations as the operators $\sigma_{xa}$. On the other hand, it follows that $\operatorname{tr}(\tilde{\sigma}_{x}B)=\operatorname{tr}(\tilde{\sigma}_{x^{\prime}}B)$ for every $B\in\mathscr{B}$, because we know that $\operatorname{tr}(\sigma_{x}B)=\operatorname{tr}(\sigma_{x^{\prime}}B)$ by assumption. Because the operators $\tilde{\sigma}_{x},\tilde{\sigma}_{x^{\prime}}$ are themselves elements in $\mathscr{B}$, it follows that $\tilde{\sigma}_{x}=\tilde{\sigma}_{x^{\prime}}$ for all $x,x^{\prime}\in\mathcal{I}_{A}$. Thus, $\tilde{\sigma}_{xa}$ is strongly non-signaling with respect to $\mathcal{B}(\mathcal{H})$ and produces the same correlations as the original strategy. ∎
Conversely, any quantum (spatial) strategy for the nonlocal game gives rise to one for the sequential game that satisfies the hypotheses of Proposition 5.11. Simply let $\mathcal{H}=\mathcal{H}_{B}$, $\sigma_{xa}=\operatorname{tr}_{A}((M_{xa}\otimes\mathbbm{1}_{B})\lvert\psi\rangle\langle\mathopen{}\psi\rvert)$, and use the operators $B_{yb}=N_{yb}$. Then strong non-signaling is satisfied with $\mathscr{B}=\mathcal{B}(\mathcal{H})$. We summarize:
@@
  • (i)
    The quantum spatial correlation set $C_{qs}$ consists precisely of the correlations produced by quantum sequential strategies that are strongly non-signaling with respect to $\mathcal{B}(\mathcal{H})$.
  • (ii)
    The quantum correlation set $C_{q}$ consists precisely of the correlations produced by finite-dimensional quantum sequential strategies that are strongly non-signaling (we can but need not assume that $\mathscr{B}=\mathcal{B}(\mathcal{H}$)).
The preceding Corollary 5.13 only characterizes the correlations produced by strongly non-signaling quantum correlations in special cases. Indeed, Definition 5.9 only requires equality on the subalgebra generated by Bob’s measurement operators. As such, the preceding results are instructive for building intuition but are insufficient for our main application, which requires the general infinite-dimensional case. Here, unlike in finite dimensions, it is not possible to reduce to the case that $\mathscr{B}=\mathcal{B}(\mathcal{H})$, and indeed we will find that the strongly non-signaling condition corresponds to general commuting operator correlations, as we prove in the next section.
5.4 A characterization of quantum commuting correlations
In this subsection, we find that the strong non-signaling condition precisely characterizes the commuting operator correlations in the sense of Definition 3.5. To establish this result, it is useful to define the following equivalent $C^{*}$-algebraic model.